OpenStudy (anonymous):

Which of the following would be equivalent to 32 ⋅ 35? 3 to the 10th power over 3 to the 3rd power 3 to the 10th power over 3 to the 7th power 31 ⋅ 37 30

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

can u help ill give more medals and i will really appreciate it

OpenStudy (anonymous):

is this 3^2 x 3^5? this simply adds the power... it should be 3^(2+5) = 3^7....

OpenStudy (anonymous):

im not exactly sure im given this exacty as u see it and told to complet it

OpenStudy (anonymous):

for any number x raised to a power of "a" and "b"... if we go to division such as... x^a over x^b = x^(a-b)... the first two expression might have the same answer as 3^7....

OpenStudy (anonymous):

can you see it?

OpenStudy (anonymous):

yeah it would Be (b) correct?

OpenStudy (anonymous):

could it be the first one? 3^10/3^3 = 3^(10-3) = 3^7.... because the 2nd one is 3^10/3^7 = 3^(10-7) = 3^3 not equal to 3^7...:-)

OpenStudy (anonymous):

yeah that makes more sense aha thanks u do a better job then my teacher

OpenStudy (anonymous):

12 to the 16th power over 12 to the 4th power. how would i simplify this

OpenStudy (anonymous):

12^16/12^4 = 12^(16-4)=?

OpenStudy (anonymous):

144

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

no it should 12^12 = 12 x 12 x 12 x 12 x 12 x 12 x 12 x 12 x 12 x 12 x 12 x 12 = 8,916,100,448,256 or approx 8.9 trillion... and not equal to 144 = 12^2 = 12 x 12

OpenStudy (anonymous):

just remember the following rules for powers or exponents: \[x^a \times x^b = x^{a+b}\]\[x^a / x^b = x^{a-b}\]\[(x^a)^b = x^{ab}\]\[(x^a)^{\frac{1}{b}} = x^{\frac{a}{b}}=\sqrt[b]{x^a} \]

OpenStudy (anonymous):

thanks @Orion1213